A candy mix that sells for $2.50 a pound is to be mixed with a candy mix that sells for $5.00 a pound. How much of each should be used to make 50 pounds of mix that sells for $4.00 a pound?
Answer by mananth(16949)

Price ----------------quantity
Candy mix I 2.5----------------x pounds
Candy mix II 5 ------ 50-xpounds
Mixture 4 ---------------- 50
2.50x+5(50-x)=4 *50
250x+500(50-x)= 20000
250x+25000-500x =20000
250x-500x=20000 -25000
-250 x = -5000
/ -250
x= 20pounds Candy mix I
30 pounds Candy mix II

m.ananth@hotmail.ca